# Ulrich complexity and Picard rank of bidouble Planes

### v1

@@ Abstract.
We determine the Picard number and the Ulrich complexity of general bidouble covers of the projective plane, providing the first systematic study of Ulrich bundles on non-cyclic abelian covers. For a bidouble plane branched along three smooth curves of degrees $n_{1},n_{2},n_{3}$, we show that $\rho(S)=1$ unless $(n_{1},n_{2},n_{3})$ belongs to an explicit list, thereby extending Buium’s classical results on double planes to the non-cyclic case. As an application, we determine the range of branch degrees for which Ulrich line bundles could exist, and we show that every even bidouble plane carries a special rank two Ulrich bundle. Our method combines the invariant-theoretic decomposition of $H^{2}(S,{\mathbb{Q}})$ under the Galois group with cohomological criteria for Ulrich bundles.
@@ 1. Introduction
The computation relies on the eigenspace decomposition of $H^{2}(S,\mathbb{Q})$ under the $(\mathbb{Z}/2)^{2}$-action and on a detailed analysis of the Hodge structure associated to the intermediate double covers determined by the three involutions of $S$. One consequence is that, for general choices of branch curves, the surface $S$ carries no Ulrich line bundles with respect to $\mathscr{O}_{S}(1)$: the Néron–Severi group is too small for such a bundle to exist (cf. Theorem 4.5). We therefore turn to the existence of rank two Ulrich bundles on $S$. The smallest integer $r\geq 1$ such that there exists a rank $r$ Ulrich bundle on $S$ with respect to $H$ is called the Ulrich complexity and is denoted by $\operatorname{uc}(S,H)$. Our second main theorem establishes the Ulrich complexity of even bidouble planes.
@@
    If $(n_{1},n_{2},n_{3})\notin T_{1}\cup T_{2}$, then $\operatorname{uc}(S,\mathscr{O}_{S}(1))=2$.
@@
The construction of Ulrich rank-two vector bundles to prove Theorem 1.2 uses the Cayley–Bacharach–Serre correspondence method (cf. Theorem 5.1). Additionally, we compute the Ulrich complexity in some low-degree cases, completing the picture for even bidouble planes (cf. Proposition 4.6).
In [4, Theorem 1.1], Basu and Pal establish the existence of rank two Ulrich bundles on certain surfaces of general type. However, our results lie outside the scope of their framework, as they assume that the surfaces are polarized by their canonical bundle. This is not our situation, since we investigate Ulrich bundles with respect to a non–very ample divisor. Moreover, [4, Theorem 1.1] requires that the canonical class generate the Picard group.
The paper is organized as follows. In Section 2, we recall the basic facts about Ulrich bundles and bidouble covers. In Section 3, we study the Nerón–Severi group of bidouble planes: first, we show that the Picard number of a bidouble plane is determined by those of its intermediate covers; second, we establish the numerical conditions under which the Picard number is strictly larger than one, leading to the proof of Theorem 1.1. In Section 4, we show that the existence of an Ulrich line bundle forces the Picard number to be greater than one. For double planes, it was shown in [22] that the only obstruction is having Picard number one. In contrast, we exhibit families of bidouble covers with Picard number greater than one that nevertheless admit no Ulrich line bundles. On the positive side, we also construct explicit Ulrich line bundles in the cases $(n_{1},n_{2},n_{3})=(0,2,2)$ and $(0,2,4)$. Finally, in Section 5, we prove Theorem 1.2, showing that the Ulrich complexity of an even bidouble plane with sufficiently large degree is equal to $2$.
@@ 2.1. Ulrich bundles
Definition 2.4.
Let $X$ be a smooth projective variety of dimension $d$, and let $H$ be an ample and globally generated divisor on $X$. A rank $2$ vector bundle $\mathscr{E}$ on $X$ is called special Ulrich if it is Ulrich with respect to $H$ and $c_{1}(\mathscr{E})=(d+1)H+K_{X}$.
Corollary 2.5 ([12, Proposition 5.1.3]).
Let $(S,H)$ be a smooth projective polarized surface and let $\mathscr{E}$ be a rank $2$ vector bundle on $S$. Then $\mathscr{E}$ is a special Ulrich bundle if and only if it satisfies
  1. (1)
    $c_{1}(\mathscr{E})\equiv 3H+K_{S},$
  2. (2)
    $c_{2}(\mathscr{E})=\frac{1}{2}(5H^{2}+3H\cdot K_{S})+2\chi(\mathscr{O}_{S}),$
  3. (3)
    $h^{0}(S,\mathscr{E})>0,$ and
  4. (4)
    $h^{0}(S,\mathscr{E}(-H))=0.$
A vector bundle is called initialized with respect to $H$ if it satisfies the conditions (3) and (4).
A fruitful approach for constructing rank 2 vector bundles on surfaces with prescribed Chern classes is provided by Cayley–Bacharach–type results. For this reason, we state the following theorem, which will play a central role in the proof of Theorem 1.2.
Theorem 2.6 ([26, Theorem 7]).
Let $X$ be a smooth, projective variety of dimension at least $2$, and let $Z\subset X$ be a subscheme of pure codimension $2$. The following statements are equivalent:
  1. (1)
    $Z$ is the zero subscheme of a section of a rank $2$ vector bundle $\mathscr{E}$,
  2. (2)
    there exist three hypersurfaces $E_{1},\,E_{2}$ and $E_{3}$ such that $E_{1}$ and $E_{2}$ have no common components, $Z=E_{1}\cap E_{2}\setminus(E_{1}\cap E_{2}\cap E_{3})$, and $E_{1}\cap E_{2}\cap E_{3}$ is a pure codimension $2$ and Cohen–Maculay.
If (1) or (2) holds, then $\mathscr{E}$ fits into a short exact sequence
$0\longrightarrow\mathscr{O}_{X}\longrightarrow\mathscr{E}\longrightarrow\mathscr{I}_{Z}(E_{1}+E_{2}-E_{3})\longrightarrow 0,$ (3)
which implies $c_{1}(\mathscr{E})=E_{1}+E_{2}-E_{3}$ and $c_{2}({\mathscr{E}})=\#Z$.
Remark 2.7.
The sequence (3) can be found at the end of the proof of [26, Theorem 7].
@@ 4.2. Ulrich line bundles in low degree
5. rank two Ulrich bundles on general even bidouble planes
In this section, we prove Theorem 1.2. We show that, in contrast to the situation for Ulrich line bundles, special Ulrich bundles exist for almost every even generic bidouble cover $S$ of the projective plane. Recall that $H$ denotes the polarization $\pi^{*}\mathscr{O}_{{\mathbb{P}}^{2}}(1)$.
By Corollary 2.5, in order to construct a special Ulrich bundle with respect to $H$, it suffices to produce a rank-two vector bundle $\mathscr{E}$ satisfying
$c_{1}(\mathscr{E})\equiv 3H+K_{S},\qquad c_{2}(\mathscr{E})=\frac{1}{2}(5H^{2}+3H\cdot K_{S})+2\chi(\mathscr{O}_{S}),$
together with the cohomological conditions
$h^{0}(S,\mathscr{E})>0,\qquad h^{0}(S,\mathscr{E}(-H))=0.$
The key idea is to apply Theorem 2.6 to construct a vector bundle of rank two satisfying these requirements. More precisely, we choose hypersurfaces $E_{1}$, $E_{2}$, $E_{3}$ such that
$3H+K_{S}\equiv E_{1}+E_{2}-E_{3},\qquad\text{and}\qquad\#Z=\frac{1}{2}(5H^{2}+3H\cdot K_{S})+2\chi(\mathscr{O}_{S}),$
where $Z=(E_{1}\cap E_{2})\setminus(E_{1}\cap E_{2}\cap E_{3})$. According to Theorem 2.6, these data produce a rank-two vector bundle $\mathscr{E}$ with the desired Chern classes. The remainder of the proof consists in using the short exact sequence (3) to verify the required cohomology vanishings.
Theorem 5.1 (Ulrich bundles of rank two).
Let $\pi\colon S\to\mathbb{P}^{2}$ be an even smooth bidouble plane branched along smooth curves $D_{1},D_{2},D_{3}$ of degrees $n_{1},n_{2},n_{3}$, respectively. If no permutation of $(n_{1},n_{2},n_{3})$ is equal to $(0,2,2)$, then $S$ carries a special Ulrich bundle $\mathscr{E}$ with respect to $H$.
Proof.
Let $_{i}=n_{i}/2, m=m$$m_{1}+m_{2}+m_{3}$ and $M=m^{2}+m_{1}^{2}+m_{2}^{2}+m_{3}^{2}$, and notice that $M$ is even. By Corollary 2.5, a rank-two vector bundle $\mathscr{E}$ is a special Ulrich bundle with respect to $H$ if and only if it satisfies
$c_{1}(\mathscr{E})\equiv mH,\qquad c_{2}(\mathscr{E})=M,\qquad h^{0}(S,\mathscr{E})>0,\qquad\text{and}\qquad h^{0}(S,\mathscr{E}(-H))=0.$
Now, we define the hypersurfaces $E_{1}$, $E_{2}$, and $E_{3}$ according to the reduction of $M$ modulo $4$.
If $M\equiv 0\pmod{4}$, set $E_{1}\colon=\pi^{*}(\ell)$, $E_{2}\colon=\pi^{*}(C)$, and $E_{3}\colon=\pi^{*}(C^{\prime})$, where $\ell$ is a general line, and $C$ and $C^{\prime}$ are smooth plane curves in general position of degrees $M/4$, and M/4+1mM/4+1-m, respectively.
On the other hand, if $M\equiv 2\pmod{4}$, set $E_{2}\colon=\pi^{*}(C)$, and $E_{3}\colon=\pi^{*}(C^{\prime})$, where $C$ and $C^{\prime}$ are smooth plane curves in general position of degrees $(M+2)/4$, and (M+2)/4+1m(M+2)/4+1-m, respectively, and $E_{1}\colon=\pi^{*}(\ell)$, where $\ell$ is a line that intersects exactly one point of $C$ and the $D_{i}$’s.
In any case, since $m\geq 3$, $\deg(C)\geq m$, in particular, the degree of $C^{\prime}$ is positive. Due to the generality, we have that $E_{1},\,E_{2}$, and $E_{3}$ have no common point.
Let $Z\subset S$ be the zero-dimensional subscheme
$Z\coloneqq E_{1}\cap E_{2}\;\setminus\;(E_{1}\cap E_{2}\cap E_{3}).$
Since the triple intersection is empty and because of the choice of $E_{i}$’s, if $M\equiv 0\pmod{4}$, we have $Z=4\left(\frac{M}{4}\right)=M$. In the case $M\equiv 2\pmod{4}$, there is a single point in $\ell\cap C$ which if ramified via $\pi$, then $Z=4\left(\frac{M+2}{4}-1\right)+2=M$. By construction, the data $(E_{1},E_{2},E_{3},Z)$ satisfy Theorem 2.6 (2). Hence, there exists a rank two vector bundle $\mathscr{E}$ fitting into
$0\longrightarrow\mathscr{O}_{S}\longrightarrow\mathscr{E}\longrightarrow\mathscr{I}_{Z}(E_{1}+E_{2}-E_{3})\longrightarrow 0.$ (13)
Since $E_{1}+E_{2}-E_{3}\equiv mH$, we obtain $c_{1}(\mathscr{E})\equiv mH$, and $c_{2}(\mathscr{E})=\#Z=M$.
Now, we claim that $h^{0}(S,\mathscr{E}(-H))=0$. Twisting (13) by $\mathscr{O}_{S}(-H)$ gives
$0\longrightarrow\mathscr{O}_{S}(-H)\longrightarrow\mathscr{E}(-H)\longrightarrow\mathscr{I}_{Z}((m-1)H)\longrightarrow 0.$ (14)
Since $h^{0}(S,{\mathscr{O}}_{S}(-H))=0$, it is enough to prove that $h^{0}(S,\mathscr{I}_{Z}((m-1)H))=0$.
For the sake of contradiction, assume that there exists a nonzero section of $\mathscr{I}_{Z}((m-1)H)$, and let $D\in|(m-1)H|$ be the corresponding effective divisor. Then $Z\subset D$. Intersecting with $E_{1}\in|H|$, we obtain $D\cdot E_{1}=4(m-1)$. On the other hand, since $Z\subset E_{1}$ and every point of $Z$ contributes at least one to the intersection multiplicity, we have $D\cdot E_{1}\;\geq\;\#Z=M$. Because $m_{1}^{2}+m_{2}^{2}+m_{3}^{2}\geq\frac{1}{3}(m_{1}+m_{2}+m_{3})^{2}=\frac{1}{3}m^{2}$, we obtain $M\geq 4m^{2}/3$. For $m\geq 3$, this gives $M>4(m-1)=D\cdot E_{1}$, a contradiction. Therefore $h^{0}\bigl(S,\mathscr{I}_{Z}((m-1)H)\bigr)=0$, and hence $h^{0}(S,\mathscr{E}(-H))=0$.
Finally, the inclusion $\mathscr{O}_{S}\hookrightarrow\mathscr{E}$ induces a nonzero section $s\in H^{0}(S,\mathscr{E})$. Combining all the above, the vector bundle $\mathscr{E}$ satisfies
$c_{1}(\mathscr{E})\equiv mH,\qquad c_{2}(\mathscr{E})=M,\qquad h^{0}(S,\mathscr{E})>0,\qquad\text{and}\qquad h^{0}(S,\mathscr{E}(-H))=0,$
as desired. ∎
Proof of Theorem 1.2.
Let us recall the definition of $T_{1}$ and $T_{2}$:
To begin with, assume that $S$ is an odd bidouble cover. Theorem 4.5 implies that $S$ does not admit an Ulrich line bundle with respect to $H$. Consequently, $\operatorname{uc}(S,H)>1$.
Now, assume that $S$ is an even bidouble cover. First, Suppose that $(n_{1},n_{2},n_{3})\notin T_{1}\cup T_{2}$. Theorem 4.5 implies that $S$ does not admit an Ulrich line bundle with respect to $H$. On the other hand, Theorem 5.1 implies the existence of an Ulrich rank two vector bundle. Consequently, $\operatorname{uc}(S,H)=2$. Now, if $(n_{1},n_{2},n_{3})\in T_{1}$ Theorem 5.1 implies that $\operatorname{uc}(S,H)\leq 2$. Finally, if $(n_{1},n_{2},n_{3})\in T_{2}$ Proposition 4.6 implies that $S$ admits an Ulrich line bundle with respect to $H$, which yields $\operatorname{uc}(S,H)=1$. ∎

### v2

# Picard rank and Ulrich line bundles on bidouble planes

@@ Abstract.
We determine the Picard number and the Ulrich complexity of general bidouble covers of the projective plane, providing the first systematic study of Ulrich bundles on non-cyclic abelian covers. For a bidouble plane branched along three smooth curves of degrees $n_{1},n_{2},n_{3}$, we show that $\rho(S)=1$ unless $(n_{1},n_{2},n_{3})$ belongs to an explicit list, thereby extending Buium’s classical results on double planes to the non-cyclic case. As an application, we determine the range of branch degrees for which Ulrich line bundles could exist. Our method combines the invariant-theoretic decomposition of $H^{2}(S,{\mathbb{Q}})$ under the Galois group with cohomological criteria for Ulrich bundles.
@@ 1. Introduction
The computation relies on the eigenspace decomposition of $H^{2}(S,\mathbb{Q})$ under the $(\mathbb{Z}/2)^{2}$-action and on a detailed analysis of the Hodge structure associated to the intermediate double covers determined by the three involutions of $S$. One consequence is that, for general choices of branch curves, the surface $S$ carries no Ulrich line bundles with respect to $\mathscr{O}_{S}(1)$: the Néron–Severi group is too small for such a bundle to exist (cf. Theorem 4.5). The smallest integer $r\geq 1$ such that there exists a rank $r$ Ulrich bundle on $S$ with respect to $H$ is called the Ulrich complexity and is denoted by $\operatorname{uc}(S,H)$. Our second main theorem establishes lower bounds on the Ulrich complexity of bidouble planes.
@@
    If $(n_{1},n_{2},n_{3})\notin T_{1}\cup T_{2}$, then $\operatorname{uc}(S,\mathscr{O}_{S}(1))>1$.
@@
The paper is organized as follows. In Section 2, we recall the basic facts about Ulrich bundles and bidouble covers. In Section 3, we study the Nerón–Severi group of bidouble planes: first, we show that the Picard number of a bidouble plane is determined by those of its intermediate covers; second, we establish the numerical conditions under which the Picard number is strictly larger than one, leading to the proof of Theorem 1.1. In Section 4, we show that the existence of an Ulrich line bundle forces the Picard number to be greater than one. For double planes, it was shown in [22] that the only obstruction is having Picard number one. In contrast, we exhibit families of bidouble covers with Picard number greater than one that nevertheless admit no Ulrich line bundles. On the positive side, we also construct explicit Ulrich line bundles in the cases $(n_{1},n_{2},n_{3})=(0,2,2)$ and $(0,2,4)$.